OpenStudy (anonymous):

The top of a ladder slides down a vertical wall at a rate of 0.375 m/s. At the moment when the bottom of the ladder is 5 m from the wall, it slides away from the wall at a rate of 0.9 m/s. How long is the ladder?

OpenStudy (anonymous):

 does this help @Euler271

OpenStudy (anonymous):

hm. got the coordinators wrong. i'll try again

OpenStudy (anonymous):



OpenStudy (anonymous):

i dont think the OP is here anymore. but in the mean time.. confirm in your mind that \[x^2 + y^2 = L^2\] and that the derivative of this equation with respect to time, t is\[2x \frac{ dx }{ dt } + 2y \frac{ dy }{ dt } \] that we are also given dy/dt as -0.375 m/s, and also dx/dt at the moment when x = 5 m. hope this helps you solve for y at this instant. knowing x and y at this instant will allow you to solve for L

OpenStudy (anonymous):

\[2x \frac{ dx }{ dt } + 2y \frac{ dy }{ dt } = 0\] this is what the derivative of the equation is with respect to time (don't forget the = 0 part)